OpenStudy (lissiewalkin):

Calculate the pH of 0.0095M solution of acetic acid give that Ka = 1.76*10^-5. I started the question by using -log(0.0095) but I'm not sure if that's all I need to do. Any help is much appreciated!

OpenStudy (chmvijay):

Ka of acetic acid is 1.7 * 10^-5 it should be provided in the appendix of your chem textbook .....CH3COOH <===> H+ + CH3COO- I.....0.0095M...........0...........0 C.... -x......................+x..........+x E...0.0095M-x...........x............x Ka = [H+] [CH3COO-] / [CH3COOH] = x * x / (0.0095M - x) = x^2 / (0.0095M - x) = 1.7*10^-5 rearrange into quadratic equation x^2 = (0.0095M -x) * 1.7*10^-5 x^2 +(1.7*10^-5)*X -(0.0096M)*(1.7*10^-5) = 0 x = 0.00040M but [H+] = x [H+] = 0.00040M pH = -log[H+] = 3.40

OpenStudy (lissiewalkin):

ohh that makes so much more sense now! thanks again chmvijay you have been a great help! :)

OpenStudy (chmvijay):

ur welcome :)